# EarthMind: Towards Multi-Granular and Multi-Sensor Earth Observation with Large Multimodal Models

### v1

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated strong performance in various vision-language tasks. However, they often struggle to comprehensively understand Earth Observation (EO) data, which is critical for monitoring the environment and the effects of human activity on it. In this work, we present EarthMind, a novel vision-language framework for multi-granular and multi-sensor EO data understanding. EarthMind features two core components: (1) Spatial Attention Prompting (SAP), which reallocates attention within the LLM to enhance pixel-level understanding; and (2) Cross-modal Fusion, which aligns heterogeneous modalities into a shared space and adaptively reweighs tokens based on their information density for effective fusion. To facilitate multi-sensor fusion evaluation, we propose EarthMind-Bench, a comprehensive benchmark with over 2,000 human-annotated multi-sensor image-question pairs, covering a wide range of perception and reasoning tasks. Extensive experiments demonstrate the effectiveness of EarthMind. It achieves state-of-the-art performance on EarthMind-Bench, surpassing GPT-4o despite being only 4B in scale. Moreover, EarthMind outperforms existing methods on multiple public EO benchmarks, showcasing its potential to handle both multi-granular and multi-sensor challenges in a unified framework.
Figure 1: The proposed EarthMind supports unified multi-granular understanding for Earth Observation (EO) imagery, including image-level, region-level, and pixel-level tasks. In addition, it enables complementary multi-sensor fusion across Optical and SAR modalities.
@@ 1 Introduction
Large multimodal models (LMMs), which integrate large language models (LLMs)  [1, 2] with visual encoders, have shown remarkable success across a variety of vision-language tasks, including image captioning [3, 4], visual question answering [5, 6, 7], and grounding [8, 9, 10]. Among various application domains, Earth Observation (EO) [11, 12, 13] is of particular importance, as it allows monitoring the Earth and the effect of human activities on it. However, LMMs trained on general-purpose images often struggle to generalize to EO data, due to a significant domain gap. Recent work has addressed this challenge by constructing large-scale instruction tuning datasets [14, 15, 16, 17, 18] specifically tailored for EO, enabling better adaptation of LMMs to this domain.
Despite recent advances, existing LMMs are limited in their understanding of EO data. Firstly, EO tasks span multiple levels of granularity, from pixel-level segmentation [19, 20], over region-level semantic understanding [21, 22], up to image-level scene classification [23, 24]. Secondly, EO data comprise multiple sensing modalities, including optical imagery (e.g., RGB and Multispectral) and Synthetic Aperture Radar (SAR) [25, 26]. These modalities are inherently complementary: optical images provide rich texture and spectral information under favorable conditions, while SAR captures structural details regardless of weather or illumination. Although various sensor types exist, effective fusion, particularly between SAR and optical data modalities, remains a key challenge for EO understanding. As summarized in Tab. 1, achieving fine-grained, multi-sensor comprehension in EO remains largely unresolved.
To tackle these challenges, we introduce EarthMind, the first LMM capable of fusing multi-sensor EO inputs and performing reasoning across multiple semantic levels, as shown in Fig. 1. It achieves this by projecting heterogeneous features from different sensors and scales into a unified semantic space, thus enabling effective interpretation by LLMs. The novelty of EarthMind lies in two key design concepts that enable spatial grounding and cross-modal understanding in EO data. First, Spatial Attention Prompting (SAP) enhances pixel-level grounding by explicitly extracting and reallocating attention to regions aligned with queried objects. This overcomes limitations of prior approaches [9, 10] that combine segmentation foundation models [27, 28] with LLMs but degrade in EO settings due to vague boundaries and scale imbalances. Second, a Cross-modal Fusion mechanism, built upon token-level contrastive learning, guides the integration of complementary modalities (e.g., RGB and SAR) into a unified semantic space. Equipped with Modality Mutual Attention, EarthMind adaptively selects the most informative features from each modality, thereby facilitating robust autoregressive learning within the LLM.
Additionally, we propose EarthMind-Bench, a new benchmark designed to evaluate LMMs in challenging EO scenarios. As shown in Tab. 1, EarthMind-Bench offers several unique features. Firstly, it encompasses multi-granular tasks, ranging from coarse-grained image understanding to fine-grained segmentation. Secondly, it introduces multi-sensor data, in particular paired RGB-SAR imagery, enabling evaluation of cross-modal fusion capabilities. Thirdly, it covers multi-level questions, spanning low-level perception as well as high-level reasoning. EarthMind-Bench consists of more than 2,000 multiple-choice and open-ended questions, providing a comprehensive benchmark to assess the ability of LMMs to interpret and reason over EO data.
We implemented EarthMind based on Qwen-2.5-3B [29], and its effectiveness is demonstrated from three key perspectives. Firstly, it achieves state-of-the-art performance on several downstream tasks, outperforming existing EO-focused LMMs of larger scale, and surpassing specialized segmentation models for the first time on pixel-level tasks. Secondly, it significantly outperforms the baselines on EarthMind-Bench across both RGB and SAR, where it improves performance through the effective fusion mechanism.
Table 1: (Left) Comparison of EarthMind with existing EO LMMs. EarthMind supports both multi-granular and multi-sensor understanding. (Right) Comparison of EarthMind-Bench with existing EO benchmarks in terms of multi-sensor support, granularity, and task level. “S” denotes single-modality input; “M” indicates multiple modalities (used independently); “F” represents paired sensor fusion. “MC” and “OE” refer to multiple-choice and open-ended formats, respectively.
Method Multi-Granular Multi-Sensor Image Region Pixel Handling Fusion RSGPT [18] GeoChat [14] EarthGPT [17] Earthmarker [30] LHRS-bot [15] SkyEyeGPT [16] Skysensegpt [16] EarthDial [31] GeoPixel [32] RSUniVLM [33] EarthMind
Benchmark Multi Sensor Multi Gran. Multi Level Task Type RSIEval [18] S MC + OE HnstD [34] S MC + OE GeoChat-Bench [14] S OE VRSBench [35] S MC + OE LHRS-Bench [15] S MC VLEO-Bench [36] S MC + OE FIT-RSRC [37] S MC UrBench [38] S MC XLRS-Bench [39] S MC + OE GEOBench-VLM [40] M MC + OE EarthMind-Bench M + F MC + OE
@@ 2 Related Work
Earth Observation LMMs. Starting from the excellent foundation of image LMMs, many works [18, 14, 16, 17, 30, 15, 37, 32, 31] try to transfer the success of general image understanding to the EO field. A key challenge is the lack of instruction-tuned EO datasets. To address this, RSGPT [18] proposed the first large-scale EO image-text paired dataset, enabling conversation tasks such as image captioning and VQA. GeoChat [14] and SkyEyeGPT [16] extend LMM capabilities to region-level visual grounding by introducing region-centric instruction data. LHRS-Bot [15] leverages large-scale available EO imagery aligned with OpenStreetMap annotations to improve VLM pretraining. To enhance complex reasoning, SkysenseGPT [37] introduces the FIT-RS dataset that focuses on understanding the relationships between spatial entities. GeoPixel [32] further pushes the boundary to pixel-level grounding by constructing an automatic pipeline to generate grounded conversations based on EO imagery. Beyond optical data, EarthDial [31] incorporates a diverse set of modalities, including Multispectral and Hyperspectral imagery and SAR, with the goal of improving generalization across heterogeneous EO data sources. Despite these advances, current EO LMMs remain limited in jointly supporting multi-granular tasks and multi-sensor fusion, which are essential for real applications.
@@
Earth Observation Multimodal Benchmarks. The rapid progress of LMMs in the EO domain has also stimulated the development of dedicated evaluation benchmarks. RSIEval [18] provides human-annotated captions and VQA pairs to evaluate VLMs in remote sensing. LHRS-Bench [15] introduces hierarchical taxonomies to assess LMMs across multiple dimensions. VLEO-Bench [36] focuses on real-world applications, including urban monitoring, disaster relief, and land use. Beyond conversational tasks, VRSBench [35] and GeoChat-Bench [14] incorporate grounding tasks to evaluate the localization capabilities of LMMs. HnstD [34] aims to detect model hallucinations, while FIT-RSRC [37] targets reasoning about object relationships. XLRS-Bench [39] leverages ultra-high-resolution imagery for comprehensive evaluation. Concurrently, GEOBench-VLM [40] proposes a multi-task geospatial benchmark covering diverse EO tasks. Despite these advances, a notable gap remains: None of the existing benchmarks explicitly evaluates the ability of LMMs to perform multi-sensor fusion, which is a key capability in real EO scenarios.
3 EarthMind
3.1 Overview
From dense urban mapping to natural terrain analysis, EO applications require models that can unify multi-modal data to enable rich scene characterization and improved decision-making. At its core, EarthMind integrates a set of vision encoders with a language model to enable flexible and generalizable cross-modal reasoning.
As shown in Fig. 2, for spatial and semantic understanding at different levels of granularity, our model employs three specialized encoders: a visual encoder $\mathbf{E}_{v}$ for global semantic perception, a region encoder $\mathbf{E}_{r}$ for object-level understanding, and a grounding encoder $\mathbf{E}_{g}$ for fine-grained spatial segmentation. These components produce hierarchical representations of the input, which are projected into a shared language space using a vision-language projector (VLP). This yields a token sequence $X^{V}=\{x^{v}_{1},x^{v}_{2},...,x^{v}_{P}\}$ that captures both global and local visual information.
To further enable dense prediction tasks, we introduce a set of learnable segmentation tokens $X^{S}=\{x^{s}_{1},x^{s}_{2},...,x^{s}_{Q}\}$, where each token can be interpreted as a query embedding tailored to capture a specific aspect of the spatial layout. The combined sequence of visual tokens $X^{V}$, segmentation tokens $X^{S}$, and tokenized language queries is processed by an LLM, which performs joint cross-modal reasoning. The hidden states of “<SEG>” tokens and the output of $\mathbf{E}_{g}$ are passed to a lightweight mask decoder $\mathbf{D}_{m}$ to produce fine-grained segmentation maps. To further improve spatial grounding in complex EO scenes, we propose Spatial Attention Prompting (SAP), detailed in Sec. 3.2.
@@ 3.1 Overview
In parallel, our framework supports multi-sensor fusion by treating heterogeneous EO data as a form of video-like input. Inspired by recent advances in video-language models [41, 42], we adopt a unified data formatting strategy: single- and dual-channel SAR images are padded to form pseudo-RGB frames, while multispectral bands are grouped in triplets to construct temporally-ordered multi-frame sequences. These sequences mimic the temporal structure of video inputs and are processed by shared encoders, enabling our model to exploit cross-frame dependencies and spectral complementarity.
To enhance inter-modal alignment, we propose a Cross-modal Fusion module (details in Sec. 3.3) that leverages paired sensor inputs under a unified language conditioning. This module facilitates effective interaction between modalities by attending to complementary cues across spectral and spatial dimensions.
Figure 2: An overview of the EarthMind, which can handle multi-granular and multi-sensor EO data understanding. It facilitates effective interaction between modalities with Cross-modal Fusion and performs accurate pixel-level understanding with Spatial Attention Prompting.
3.2 Spatial Attention Prompting (SAP)
In the above architecture, the “<SEG>” token acts as an implicit language query that searches for potential target objects within the image. However, empirical results show a significant performance drop when applying this framework to EO scenarios. To investigate the cause of this degradation, we examine whether the “<SEG>” token effectively captures the semantics of the queried objects aligned with the query in EO imagery.
During LLM inference, tokens interact with each other through multi-head attention layers. A token that receives higher attention weights from others is considered more influential in the model’s reasoning process. Motivated by this, we measure the importance of each image token with respect to the “<SEG>” token by computing the cross-attention maps between segmentation tokens and all image tokens across all transformer layers, resulting in $A_{\text{seg}}\in\mathbb{R}^{L\times H\times Q\times P}$, where $L$ is the number of transformer layers and $H$ is the number of attention heads. We compute the attention map by applying softmax over the image token dimension, and average over the heads to obtain a Seg-to-Image attention map:
$\small\hat{A}_{\text{seg}}=\frac{1}{H}\sum_{h=1}^{H}\text{Softmax}(A_{\text{seg}}^{(l,h)}),\quad\hat{A}_{\text{seg}}\in\mathbb{R}^{L\times Q\times P}$ (1)
As shown in Fig. 2, the attention map reflects the spatial focus of the LLM’s segmentation prompts. However, EO imagery often exhibits extreme object scale variations, ambiguous textures, and weak semantic boundaries, which lead to attention drifting away from the target regions. Our attention visualizations further illustrate this phenomenon in Sec. 5.4.
To guide the model toward more semantically meaningful spatial priors, we introduce a supervision signal based on ground-truth masks. Specifically, we downsample the binary mask to match the image token resolution and treat it as a probability distribution over tokens. Then, we minimize the Kullback–Leibler (KL) divergence [43] between the normalized attention map and the target distribution:
$\small\mathcal{L}_{\text{KL}}=\sum_{l=1}^{L}\sum_{q=1}^{Q}\text{KL}\left(\hat{A}_{\text{seg}}^{(l,q)}\,\|\,G_{q}\right)$ (2)
where $G_{q}\in\mathbb{R}^{P}$ is the ground-truth token-level distribution corresponding to the $q$-th seg token. This training objective encourages the segmentation tokens to attend more accurately to the queried regions, thereby enhancing pixel-level mask prediction.
3.3 Cross-modal Fusion
Given multi-sensor input, we extract cross-modal image tokens before feeding into the LLM: $X^{V}_{o}=\{x^{o}_{1},x^{o}_{2},...,x^{o}_{P}\}$ and $X^{V}_{s}=\{x^{s}_{1},x^{s}_{2},...,x^{s}_{P}\}$, which represent features from optical (RGB) and non-optical (e.g., SAR) modalities, respectively. To fuse modalities, we propose the concepts of Modality Alignment and Modality Mutual Attention, as detailed below.
Modality Alignment. To facilitate effective multi-sensor fusion, we adopt an online contrastive learning strategy to align non-optical features (e.g., SAR) with the optical (RGB) feature space, leveraging the pretrained vision-language projector for optical inputs. Specifically, given a training batch of size $B$, we define the cross-modal contrastive loss as:
$\small\mathcal{L}_{CL}=-\frac{1}{B}\sum_{i=1}^{B}\left(\log\frac{f(x^{o}_{i},x^{s}_{i})}{f(x^{o}_{i},x^{s}_{i})+\sum\limits_{k\neq i}f(x^{o}_{i},x^{s}_{k})}+\log\frac{f(x^{s}_{i},x^{o}_{i})}{f(x^{s}_{i},x^{o}_{i})+\sum\limits_{k\neq i}f(x^{s}_{i},x^{o}_{k})}\right)$ (3)
where $f(\cdot)$ represents the cross-modal cosine similarity operation between paired tokens.
Modality Mutual Attention. After aligning different modalities into a shared semantic space, we further enable the model to dynamically assess the information content of each modality, thereby emphasizing the most informative representations for downstream LLM reasoning.
Specifically, we first introduce a set of learnable queries $Q\in\mathbb{R}^{k\times D}$ to extract neighborhood-aware features from each modality. For example, the representation of SAR is computed via element-wise multiplication: $\hat{X}^{V}_{s}=Q\odot X^{V}_{s}$, where $\hat{X}^{V}_{s}\in\mathbb{R}^{k\times P\times D}$. RGB features are processed similarly to obtain $\hat{X}^{V}_{o}$. We then compute cross-modal importance weights using a dot product:
$\small W_{s}=\text{Softmax}(X^{V}_{o}(\hat{X}^{V}_{s})^{\top}),\quad W_{o}=\text{Softmax}(X^{V}_{s}(\hat{X}^{V}_{o})^{\top})$ (4)
where the weights $W_{s},W_{o}\in\mathbb{R}^{P}$ reflect the relevance of the SAR and RGB tokens in cross-modal contexts, by which we reweight $X^{V}_{s}$ and $X^{V}_{o}$ accordingly and concatenate them into the LLM for joint understanding.

### v2

# EarthMind: Leveraging Cross-Sensor Data for Advanced Earth Observation Interpretation with a Unified Multimodal LLM

@@ Abstract
Earth Observation (EO) data analysis is vital for monitoring environmental and human dynamics. Recent Multimodal Large Language Models (MLLMs) show potential in EO understanding but remain restricted to single-sensor inputs, overlooking the complementarity across heterogeneous modalities. We propose EarthMind, a unified vision-language framework that handles both single- and cross-sensor inputs via an innovative hierarchical cross-modal attention (i.e., HCA) design. Specifically, HCA hierarchically captures visual relationships across sensors and aligns them with language queries, enabling adaptive fusion of optical and Synthetic Aperture Radar (SAR) features. To support cross-sensor learning, we curate FusionEO, a 30K-pair dataset with diverse annotations, and establish EarthMind-Bench, a 2,841-pair benchmark with expert annotations for perception and reasoning tasks. Extensive experiments show that EarthMind achieves state-of-the-art results on EarthMind-Bench and surpasses existing MLLMs on multiple EO benchmarks.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have revolutionized vision-language understanding by integrating powerful language models (OpenAI, 2023; Team et al., 2023) with visual encoders, achieving remarkable performance across diverse tasks, including image captioning (Xu et al., 2024; Chen et al., 2024a), visual question answering (Li et al., 2023; Liu et al., 2023a; Li et al., 2024a), and visual grounding (Lai et al., 2024; Rasheed et al., 2024; Yuan et al., 2025). Earth Observation (EO) represents a particularly critical application domain with far-reaching implications for environmental monitoring (Wójtowicz et al., 2016) and disaster response (Van Westen, 2000). However, the distinctive characteristics of EO imagery create a substantial domain gap that limits the effectiveness of MLLMs trained on natural images. Recent efforts have addressed this challenge by developing EO-specific instruction tuning datasets (Kuckreja et al., 2024; Muhtar et al., 2024; Zhan et al., 2025; Zhang et al., 2024b; Hu et al., 2025) and specialized architectures, demonstrating significant improvements in adapting MLLMs to remote sensing applications.
Despite these advances, existing MLLMs for EO remain fundamentally limited by their reliance on single-sensor inputs, failing to exploit the complementary nature of heterogeneous sensing modalities (Fuller et al., 2023; Liu et al., 2024a). Modern EO platforms routinely provide synchronized multimodal data: optical sensors like Sentinel-2 capture high-resolution multispectral imagery rich in spectral and textural information, while SAR sensors like Sentinel-1 deliver all-weather, day-night observation capability. These modalities exhibit natural complementarity—optical sensors excel in clear conditions with detailed spectral signatures but suffer from cloud occlusion and illumination dependency, whereas SAR penetrates atmospheric interference but exhibits lower signal-to-noise ratios and speckle noise. As illustrated in Fig. 1, this fundamental trade-off underscores that no single sensor can comprehensively capture scene information, necessitating advanced and effective cross-sensor fusion strategies for robust and comprehensive EO data understanding.
However, achieving effective cross-sensor fusion presents challenges. The fundamental heterogeneity between optical and SAR data, rooted in distinct imaging principles and signal characteristics, makes naive approaches like token concatenation ineffective. While traditional fusion methods (Zhang et al., 2025a; Schmitt et al., 2017; Li et al., 2022) show promise in specific contexts, they remain constrained by task-specific architectures and lack the flexibility for diverse vision-language tasks.
To address these limitations, we present EarthMind, a unified MLLM tailored for EO that seamlessly integrates both single-sensor and cross-sensor data. This integration enhances representational capacity, enabling the model to couple global scene context with local spatial detail and thereby supporting tasks ranging from high-level reasoning to fine-grained pixel-level grounding. At the heart of EarthMind is the hierarchical cross-modal attention (i.e., HCA) design, a lightweight mechanism that equips MLLMs with adaptive cross-sensor fusion. By learning spatially adaptive weights tied to sensor density and query relevance, HCA mitigates modality imbalance and preserves complementary SAR cues otherwise suppressed by optical bias. To overcome the scarcity of paired cross-sensor data, we further introduce FusionEO, a 30K-sample instruction-tuning dataset generated through an automated pipeline, ensuring the semantic diversity and task coverage required for effective training.
Additionally, we introduce EarthMind-Bench, a large-scale benchmark for evaluating cross-sensor MLLMs in EO scenarios. It enables flexible evaluation with both single- and multi-sensor inputs, spans tasks from image-level understanding to pixel-level segmentation, and covers reasoning from perception to high-level analysis. Comprising 2,841 expertly annotated pairs, EarthMind-Bench provides a unified platform to rigorously assess MLLM capabilities in interpreting and reasoning over EO data. Overall, the effectiveness of EarthMind is demonstrated in two key aspects:
  • It successfully exploits cross-sensor complementarity to enhance diverse EO understanding tasks—including dense captioning, VQA, and segmentation—achieving state-of-the-art performance on our multi-sensor EarthMind-Bench with only 4B parameters.
  • It also demonstrates superior performance on single-sensor benchmarks across various tasks, spanning image-level classification and VQA, region-level visual grounding, and pixel-level referring expression segmentation.
Figure 1: From a single Optical or SAR image, GeoChat (Kuckreja et al., 2024) and GeoPixel (Shabbir et al., 2025) give inaccurate VQA or segmentation results. EarthMind first utilize cross-sensor information to boost several tasks.
@@ 2 Related Work
Earth Observation MLLMs. Building upon the success of general image MLLMs, numerous works (Hu et al., 2025; Kuckreja et al., 2024; Zhan et al., 2025; Zhang et al., 2024b; a; Muhtar et al., 2024; Luo et al., 2024; Shabbir et al., 2025; Soni et al., 2024) have attempted to transfer these capabilities to the EO domain. A key challenge is the scarcity of instruction-tuned EO datasets. RSGPT (Hu et al., 2025) addressed this by proposing the first large-scale EO image-text paired dataset, enabling conversational tasks such as image captioning and VQA. GeoChat (Kuckreja et al., 2024) and SkyEyeGPT (Zhan et al., 2025) extended capabilities to region-level visual grounding through region-centric instruction data. LHRS-Bot (Muhtar et al., 2024) leverages large-scale EO imagery aligned with OpenStreetMap annotations to improve pretraining. To enhance complex reasoning, SkysenseGPT (Luo et al., 2024) introduces the FIT-RS dataset focusing on spatial entity relationships. GeoPixel (Shabbir et al., 2025) pushes the boundary to pixel-level grounding through an automated pipeline for generating grounded conversations. Beyond optical data, EarthDial (Soni et al., 2024) incorporates diverse modalities including multispectral, hyperspectral, and SAR data to improve generalization. Despite these advances, current MLLMs remain limited in leveraging multi-sensor inputs for comprehensive EO understanding.
@@
Earth Observation Multimodal Benchmarks. The rapid development of EO MLLMs has stimulated dedicated evaluation benchmarks. RSIEval (Hu et al., 2025) provides human-annotated captions and VQA pairs for remote sensing evaluation. LHRS-Bench (Muhtar et al., 2024) introduces hierarchical taxonomies for multi-dimensional assessment. VLEO-Bench (Zhang & Wang, 2024) focuses on real-world applications including urban monitoring and disaster relief. Beyond conversational tasks, VRSBench (Li et al., 2024b) and GeoChat-Bench (Kuckreja et al., 2024) incorporate grounding tasks for localization evaluation. HnstD (Pang et al., 2025) targets hallucination detection, while FIT-RSRC (Luo et al., 2024) evaluates object relationship reasoning. XLRS-Bench (Wang et al., 2025) leverages ultra-high-resolution imagery, and GEOBench-VLM (Danish et al., 2024) proposes a multi-task geospatial benchmark. Despite these advances, none explicitly assess MLLM performance under multi-sensor scenarios, where complementary cues across heterogeneous modalities are critical for robust EO understanding.
3 Method
From dense urban mapping to natural terrain analysis, EO applications require models that can natively fuse heterogeneous sensors for reliable scene understanding and decision-making. To this end, we present EarthMind, a unified vision-language framework that accepts either single-sensor or paired multi-sensor inputs and produces task-appropriate outputs, including natural-language responses and pixel-accurate masks. Sec. 3.1 outlines the overall architecture, which consists of multiple visual encoders, a vision-language projector, an LLM, and a lightweight mask decoder. Sec. 3.2 introduces our core multi-sensor fusion module, Hierarchical Cross-modal Attention (HCA). Sec. 3.3 describes the paired cross-sensor instruction-tuning corpus FusionEO and training objectives.
Figure 2: An overview of EarthMind, which supports both single-sensor and multi-sensor EO data understanding, including high-level image reasoning and pixel-level grounding tasks. The framework facilitates cross-sensor fusion through Hierarchical Cross-modal Attention (HCA), which selectively combines complementary features from optical and SAR data guided by both cross-modal visual relationships and input query relevance.
3.1 EarthMind Architecture
Fig. 2 illustrates the architecture of EarthMind. Our model employs a visual encoder $\mathbf{E}_{v}$ for global semantic perception and a grounding encoder $\mathbf{E}_{g}$ for fine-grained spatial understanding. A vision-language projector (VLP) transforms visual features into a token sequence $X^{V}$ aligned with the language space. These visual tokens, combined with input text tokens, serve as input to the LLM, which generates text predictions accordingly. To enable dense prediction tasks, we introduce a special token “<SEG>” that guides the mask decoder $\mathbf{D}_{m}$ to produce fine-grained segmentation maps.
@@ 3.1 EarthMind Architecture
Our framework supports flexible visual input by treating heterogeneous EO data as video-like sequences. Inspired by recent advances in video-language models (Maaz et al., 2023; Shu et al., 2024), we adopt a unified data formatting strategy: single- and dual-channel SAR images are padded to form pseudo-RGB frames, while multispectral bands are grouped in triplets to construct multi-frame sequences. This approach mimics temporal video structure and enables shared encoder processing, allowing our model to exploit cross-frame dependencies and spectral complementarity.
3.2 Cross-sensor Fusion
Motivation. Given multi-sensor visual inputs, we extract sensor-specific image tokens: $X^{V}_{o}\in\mathbb{R}^{P\times D}$ for optical and $X^{V}_{s}\in\mathbb{R}^{N\times D}$ for SAR modalities, where $D$ denotes the LLM’s embedding dimension. The most straightforward fusion approach is concatenating these tokens into a joint representation $\hat{X}^{V}\in\mathbb{R}^{(P+N)\times D}$. However, this naive concatenation presents two critical limitations. First, the increased sequence length results in quadratic computational overhead due to self-attention mechanisms. Second, despite equal representation in the input, LLMs exhibit strong bias toward optical tokens, potentially overlooking valuable SAR information. This bias likely stems from the predominance of RGB images in pre-training data.
To quantitatively measure this modality imbalance, we introduce the Modality Attention Score (MAS):
$A_{\ell}^{m}=\frac{\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{T}_{m}}\alpha_{i,j}^{\ell}}{\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{I}}\alpha_{i,k}^{\ell}}$ (1)
where $\mathcal{I}$ denotes all visual tokens, $\mathcal{T}_{m}\subset\mathcal{I}$ represents tokens from modality $m$, and $\alpha_{i,j}^{\ell}$ denotes the self-attention weight from token $i$ to $j$ at layer $\ell$ in LLM. MAS quantifies the proportion of attention allocated to each modality.
Our empirical analysis (Sec. 5.4) reveals that standard MLLMs with concatenation fusion consistently exhibit higher $A_{\ell}^{\text{optical}}$ than $A_{\ell}^{\text{SAR}}$ across layers, confirming the modality bias. This motivates our Hierarchical Cross-modal Attention module, which explicitly models cross-sensor interactions for balanced and efficient multi-modal fusion.
Hierarchical Cross-modal Attention. We propose a hierarchical attention mechanism that captures both cross-sensor complementarity and text-guided relevance for effective multimodal fusion. Our approach operates in two stages: (1) bidirectional cross-modal attention between optical and SAR features, and (2) text-guided attention to align visual features with task requirements.
Stage 1: Cross-Sensor Attention. To ensure spatial alignment between modalities with different resolutions, we apply adaptive pooling to align optical features with SAR dimensions, yielding $\hat{X}^{V}_{o}\in\mathbb{R}^{N\times D}$. We then compute bidirectional attention maps:
$A^{o2s}=\mathrm{Softmax}\!\left(\frac{\hat{X}^{V}_{o}(X^{V}_{s})^{\top}}{\sqrt{D}}\right),\quad A^{s2o}=\mathrm{Softmax}\!\left(\frac{X^{V}_{s}(\hat{X}^{V}_{o})^{\top}}{\sqrt{D}}\right)$ (2)
where $A^{o2s}_{ij}$ captures attention from optical token $i$ to SAR token $j$, and vice versa for $A^{s2o}_{ij}$.
To derive modality-specific importance scores, we aggregate attention weights across spatial dimensions:
$\alpha^{o}_{i}=\frac{1}{N}\sum_{j=1}^{N}A^{s2o}_{ji},\quad\alpha^{s}_{i}=\frac{1}{N}\sum_{j=1}^{N}A^{o2s}_{ji}$ (3)
Here, $\alpha^{o}_{i}$ represents how much attention the $i$-th optical token receives from all SAR tokens, indicating its importance from the SAR perspective.
Stage 2: Text-Guided Attention. We incorporate task-specific guidance by computing visual token relevance to the text prompt. Given text embeddings $X^{T}\in\mathbb{R}^{L\times D}$ aligned with visual features through the projector:
$A^{o2t}=\mathrm{Softmax}\!\left(\frac{\hat{X}^{V}_{o}(X^{T})^{\top}}{\sqrt{D}}\right),\quad A^{s2t}=\mathrm{Softmax}\!\left(\frac{X^{V}_{s}(X^{T})^{\top}}{\sqrt{D}}\right)$ (4)
We aggregate these scores to obtain text-relevance weights:
$\beta^{o}_{i}=\frac{1}{L}\sum_{j=1}^{L}A^{o2t}_{ij},\quad\beta^{s}_{i}=\frac{1}{L}\sum_{j=1}^{L}A^{s2t}_{ij}$ (5)
Adaptive Fusion. The final fusion weights combine cross-modal complementarity and text relevance:
$\gamma^{o}_{i}=\lambda\cdot\alpha^{o}_{i}+(1-\lambda)\cdot\beta^{o}_{i},\quad\gamma^{s}_{i}=\lambda\cdot\alpha^{s}_{i}+(1-\lambda)\cdot\beta^{s}_{i}$ (6)
where $\lambda$ is a learnable parameter balancing the two attention mechanisms. The fused representation is computed as:
$X^{V}_{\text{fused},i}=w^{o}_{i}\cdot\hat{X}^{V}_{o,i}+w^{s}_{i}\cdot X^{V}_{s,i}$ (7)
where $w^{o}_{i},w^{s}_{i}]=\text{Softmax}([\gamma^{o}_{i},\gamma^{s}_{i}])$ are the normalized fusion weights. This produces final multimodal tokens $X^{V}_{\text{fused}}\in\mathbb{R}^{N\times D}$ for LLM input, effectively reducing sequence length while preserving cross-modal information.